If 124 people attend a concert and tickets for adults cost $3 while tickets for children cost $2 and total receipts for the concert was $316, how many of each went to the concert?
Answer #1

Let there was x adults . Therefore there was (124 - x) children.

Total cost of tickets for x adults is $ 3 * x = $ 3x

Total cost of tickets for (124 - x) children is

$ 2 * (124 -x) = $ (248 - 2x)

Therefore

3x +(248-2x) = 316

3x - 2x = 316 - 248

x = 68

Then 124 - x = 124 - 68 = 56

68 adults and 56 children went to the concert.